OpenStudy (anonymous):

help... step by step please. Solve the equation for y. Then graph the equation. 5x – 4y = 12

OpenStudy (anonymous):

5x – 4y = 12 subtract 5x from both sides

OpenStudy (anonymous):

do it

OpenStudy (anonymous):

-4y=7x ?

OpenStudy (anonymous):

idk...

OpenStudy (anonymous):

no its 4y = 12-5x

OpenStudy (anonymous):

now divide both sides by 4

OpenStudy (anonymous):

???

OpenStudy (anonymous):

dont I have to subtract 12-5x first? which gives me 7x?

OpenStudy (anonymous):

no no you have this 5x – 4y = 12 subtracting 5x from both sides we get 5x-5x-4y = 12-5x -4y = 12 - 5x

OpenStudy (anonymous):

ok.... i'm confused what next?

OpenStudy (anonymous):

now dividing both sides by -4 we get y = -3+5/4x re arranging we get y= (5/4)x -3 this the first part of your question

OpenStudy (anonymous):

is that the right answer?

OpenStudy (anonymous):

yesssssss

OpenStudy (anonymous):

now take y= (5/4)x -3 to plot this frist put x =0 in it you will get some y then put y = 0 in y= (5/4)x -3 you will get some x plot both these points and join them you are done

OpenStudy (anonymous):

me coming after some time see you soon

OpenStudy (anonymous):

what

OpenStudy (anonymous):

uhmm... could you help with another question I tried doing it but I got all mixed up

OpenStudy (anonymous):

ok post it